OpenStudy (anonymous):

Seven times a number minus the number is -48. Find the number.

OpenStudy (calculusxy):

Let's write an equation with the information that we already know. n = number that we are trying to find 7n = 7 times "n" \[7n - n = -48\]

OpenStudy (calculusxy):

Combine the like terms and then find the value of n.

OpenStudy (anonymous):

so -8/-48.? that would be 6.?

OpenStudy (calculusxy):

\[7n - n = -48\] \[6n = -48\] \[n = \frac{ -48 }{ 6 }\]

OpenStudy (calculusxy):

Remember that it is subtraction.

OpenStudy (calculusxy):

@math_girl27 Did you get your answer?

OpenStudy (anonymous):

yess thank you @calculusxy